Evaluate the integral "\displaystyle{\int_{0}^{\pi/2}\sin{(x)}\ln{\sec{(x)}}dx}"

  1. 1
  2. 0
  3. 1/2
  4. 8
1
Expert's answer
2021-09-08T08:10:01-0400



"\displaystyle{\int_{0}^{\pi/2}\sin{(x)}\ln{\sec{(x)}}dx}"


Solution


"\displaystyle{\int\sin{(x)}\ln{\sec{(x)}}dx}"


Integrate by parts "\displaystyle\int fg'=fg-\int f'g"


"f=ln(sec(x))"

"f'=tan(x)"


"g'=sin(x)"

"g=-cos(x)"


Hence from "\displaystyle\int fg'=fg-\int f'g"

"=-cos(x)ln(sec(x))-\displaystyle\int-cos(x)tan(x)dx"


Now solving ;

"\displaystyle\int-cos(x)tan(x)dx"


Apply linearity;

"=-\displaystyle\int cos(x)tan(x)dx"

Now solving;

"\displaystyle\int cos(x)tan(x)dx"


Rewrite/ Simplify using trigonometric/hyperbolic identities

"=\displaystyle\int sin(x)dx"


This is the standard integrals:


"=-cos (x)"


Plug in solved integrals;

"=-cos(x)ln(sec(x))-\displaystyle\int-cos(x)tan(x)dx"


"=(-cos(x)ln(sec(x))-cos(x)"


The problem is solved


"\displaystyle{\int\sin{(x)}\ln{\sec{(x)}}dx}"


"=[-cos(x)ln(sec(x))-cos(x)+C]"


Simplify

"=cos(x)(ln(|cos(x)|)-1)+C"


Hence

 "\displaystyle{\int_{0}^{\pi/2}\sin{(x)}\ln{\sec{(x)}}dx}=
[cos(x)(ln(|cos(x)|)-1)+C]_0^{\frac{π}{2}}"


"=[cos(\frac{π}{2})(ln(|cos(\frac{π}{2})|)-1)+C]-
[cos(0)(ln(|cos(0)|)-1)+C]= 1"